OpenStudy (anonymous):

let f(x)= 5x-3 and g(x)=2x+7. Find the following functions. Formulas: (f+g)(x)=f(x)+g(x) (f-g)(x)f(x)-g(x) (fg)(x)=(f(x))(g(x)) (f*g)(x)=f(g(x)) (g*f)(x)=g(f(x)) A. (f+g)(x) B. (f-g)(x) C. (fg)(x) D. (f/g)(x) E. (f*g)(x)

OpenStudy (anonymous):

a=7x+4 b=3x-10 c=10x^2+29x-21 d=5x-3/2x+7 e=5(2x+7)-3